# A Parallel CPU-GPU Framework for Cost-Bounded DFS with Applications to IDA* and BTS

### v1

@@ Introduction
Theory
In this section, we demonstrate that Batch IDA* is guaranteed to find an optimal solution, assuming one exists. Our proof builds on the established optimality of IDA*. Since the original AIDA* paper lacks a formal proof of optimality, we begin by showing this guarantee for AIDA* and then use this result to prove the optimality of Batch IDA*.
Lemma 1.
Given an admissible monotone cost function, iterative-deepening- A* will find a solution of least cost if one exists (Korf 1985).
Lemma 2.
Proof.
From Lemma 1, we know that IDA* using the cost function $f$ finds an optimal solution if a solution exists. Therefore, our task is to demonstrate that AIDA*, using the same cost function $f$ as IDA*, expands all the nodes that IDA* expands at each cost threshold. We know that, like IDA*, AIDA* does not increase the cost threshold until all nodes with costs below the current threshold have been expanded. Assume, for the sake of contradiction, that there exists a node $s_{missed}$ that IDA* expands but AIDA* does not. Let us also assume that the initial phase in AIDA* proceeds up to a depth of $d_{init}$. We select a path from the root to $s_{missed}$, denoted as $path$. The node $s_{missed}$ falls into one of two categories:
  1. 1.
    The length of $path$ is less than or equal to $d_{init}$.
  2. 2.
    The length of $path$ exceeds $d_{init}$.
If $s_{missed}$ falls within the first category, it must be expanded during initial phase, as this phase generates the entire tree up to $d_{init}$ levels. Therefore, in this case, $s_{missed}$ cannot be missed by AIDA*. If $s_{missed}$ falls into the second category, we can express the $path$ as $\{root,s_{1},...s_{d_{init}},...,s_{missed}\}$. The $path$ intersects the tree generated in the initial phase at node $s_{d_{init}}$. We know that, during second and third phases, AIDA* grows the subtree for all leaf nodes of the initial tree, including $s_{d_{init}}$. Since both AIDA* and IDA* use the same cost function $f$, the same cost threshold, and both employ a depth-first search strategy, AIDA* is guaranteed to expand the node $s_{missed}$ during subtree expansion of $s_{d_{init}}$. Thus, AIDA* does not miss any node that IDA* expands. This concludes the proof that AIDA* will find an optimal solution if one exists. ∎
Theorem 1.
Proof.
  1. 1.
  2. 2.
    $h_{M}(s)\leq h(s)\ \ \ \ \forall s\in S.$
@@
The memory cost of IDA* is $\mathcal{O}(d)$, where $d$ is the solution depth (Korf 1985). For AIDA*, which utilizes $n$ CB-DFS threads, the memory cost increases to $\mathcal{O}(dn)$ as it concurrently expands $n$ subtrees. In Batch IDA*’s CB-DFS, each thread manages $k$ subtrees in a stack, resulting in a memory cost of $\mathcal{O}(dnk)$. We know that $n$ and $k$ are related to hardware speed on any domain. So, these parameters are expected to be constant as hardware is constant. Additionally, while the heuristic evaluations for unexpanded nodes are stored temporarily, memory is freed after their expansion, keeping the total memory cost at $\mathcal{O}(dnk)$.
Both AIDA* and Batch IDA* require loading their respective heuristics into memory. If AIDA* uses a PDB heuristic $h$ and Batch IDA* uses a learned model $h_{M}$, the total memory cost for AIDA* is $\mathcal{O}(nd+size(h))$ and for Batch IDA*, it is $\mathcal{O}(dnk+size(h_{M}))$. While PDB heuristics typically fit into memory and thus their cost is constant and negligible, large PDBs, such as the 12-edge Rubik’s cube PDB (500 GB), exceed memory capacity, making them impractical for AIDA*. This limitation underscores the importance of compressing PDB heuristics using deep learning models (Li et al. 2022) and designing search algorithms, such as the CB-DFS used by Batch IDA* and Batch BTS, to use them efficiently.
@@
SingleGPU Batch IDA* $1$ $>$ 100 - - $16.80$ $104,869$ $225,542$
SingleGPU Batch IDA* $8$ $>$ 100 - - $3.97$ $93,610$ $201,217$
@@
Table 1: Summary results comparing learning algorithms using the same PDB heuristic.
@@
MultiGPU Batch IDA*
In most servers, multiple GPU devices are available; however, SingleGPU Batch IDA* is limited to using one GPU device. This means that even if faster CPUs are employed for the search, overall performance gains are minimal because the batch processing time remains unchanged, forcing CB-DFS threads to wait until the model $h_{M}$ is ready to take a new batch. A potential solution involves loading copies of $h_{M}$ onto GPU, allowing multiple batches to be processed simultaneously. However, this can lead to oversubscription, where the GPU lacks sufficient SMs to efficiently handle all batches, potentially increasing runtime. To address this issue, we introduce MultiGPU Batch IDA* which leverages multiple GPU devices. In this approach, $h_{M}$ is loaded onto each GPU, with a batch and a batch-processing thread assigned to each device. CB-DFS threads are distributed evenly across the GPUs, with each GPU processing the batch of nodes generated by its respective CB-DFS threads. This setup maximizes CPU utilization by increasing the number of batches while ensuring sufficient GPU resources are available to process them concurrently.
Experimental Results
We evaluate the Batch IDA* algorithm to understand its performance and how it compares to existing algorithms. Our first two experiments evaluate whether our Batch IDA* design is effective in reducing the computational overhead of neural heuristics, and how the size of the neural heuristic model impacts performance. This is followed by an analysis of the impact of different hyperparameters and hardware. Finally, we perform a comparison between classical and neural heuristics to evaluate their current performance differences.
Experimental Setup
We conduct our experiments across two domains: the 3x3 Rubik’s cube and the 4x4 sliding tile puzzle (STP). We use relatively modest computational resources in our experiments. Specifically, we utilize a server equipped with 32 AMD Ryzen Threadripper 2950X CPU cores and two NVIDIA GeForce RTX 2080 Ti GPUs running CUDA version 12.4. The Batch IDA* algorithm is implemented in C++, and Libtorch is employed for the deep learning components. Any CPU parallelization uses one thread per CPU core. For the non-learning algorithms, IDA* and AIDA*, we use $h_{\text{PDB}}$ which is the 8-corners PDB for Rubik’s cube, and the sum of 1-7 and 8-15 tile additive PDBs for STP. For the STP, we use standard benchmark instances (Korf 1985). For Rubik’s cube, we generate instances by performing a random walk of length 11 to 13 starting from the goal. All PDB heuristics and baseline implementations are sourced from HOG2111https://github.com/nathansttt/hog2/tree/PDB-refactor. The heuristic $h_{M}$ for the STP is the model introduced by Li et al. (2022). It is constructed by combining two ensemble models: the first is trained on the difference between the 1-7 tile PDB and the Manhattan distance heuristic, while the second is similarly trained using the 8-15 tile PDB. We also trained an admissible heuristic $h_{M}$ using the 8-corners PDB for the Rubik’s Cube; however, a full description is outside the scope of this paper. But, what is important is that both approaches learn classifiers instead of using regression. Because of this, the IDA* search layers are well-behaved. Learning with regression would result in a broader range of non-integer heuristic values, and thus require using BTS instead of IDA*.
Performance Evaluation
Our first experiment is designed to evaluate the effectiveness of batch operations in Batch IDA*. We designed this experiment to isolate as many variables as possible in the experiment and to isolate the quality of neural heuristic from the mechanics of Batch IDA*. We isolate the impact of neural heuristic training by using randomly initialized neural heuristics. This allows us to quickly experiment with different network topologies and sizes without having to re-train neural heuristics for every setting. While it is possible to use Batch IDA* with random heuristics, this would introduce variance in tree sizes across heuristics. Additionally, random heuristics are highly inconsistent (Felner et al. 2011), which introduces the potential for the worst-case time complexities for IDA* and A* (Helmert et al. 2019; Kaur and Sturtevant 2022). We avoid this by calculating the neural heuristic as if we were going to use it, but then replace the heuristic values with values from $h_{\text{PDB}}$ to guide the CB-DFS.
We use the first 50 instances from Korf’s benchmarks for the STP domain. On this domain, going from batch size of 1 to 800 reduces the running time from 16.80s to 0.54s, a 30$\times$ improvement. Using two GPUs and batch size 800 reduces this further to 0.44s. The variance in nodes expanded is due to a deep initial phase in Batch IDA* aimed at producing sufficient work pieces, which results in batches containing extra nodes that are not necessary to solve the problem ($f$-cost greater than the threshold). Although Batch IDA* is faster than Batch A* with respect to time per node expansion, Batch A* expands fewer nodes because it detects duplicates, which Batch IDA* does not.
@@
Our second experiment looks at the impact of model size on CB-DFS performance. We run Batch IDA* on the same domains with a single GPU and the best parameters from Table 1. We vary the model size by fixing the number of layers, but changing the number of fully connected neurons in each layer. The results of this experiment are found in Table 2. These results show that increasing the size of the model does not linearly increase the time required to evaluate the model, although larger models are more expensive to evaluate. Thus, there will be trade-offs between model quality and speed when deploying neural heuristics.
The results also show that Batch A* is less affected by model size than Batch IDA*. The main reason for this difference is that the batches in Batch IDA* are not always full, meaning the GPU isn’t being used as efficiently in Batch IDA*. This can be caused by node generations above the cost thresholds and the uneven distribution of work between subtrees. When the model requires more time to evaluate, these issues are accentuated. See the Supplementary Material for more analysis of how full the batches are in pratice.
Hyperparameter Analysis
Next, we present the impact of hyperparameters on the performance of Batch IDA*. For this experiment, we consider the batch size and the number of subtrees per thread, specified by workNum in Algorithm 3. The timing results for different combinations of hyperparameter values are shown in Figure 3. In Rubik’s Cube there is only about 10% variation between the best and worst parameters, whereas in STP there is a 1.84$\times$ difference. The STP domain has less balanced subtrees than Rubik’s Cube. With too many subtrees, the work is divided out to all threads too quickly, reducing the ability to load-balancing between smaller and larger subtrees. The range of batch sizes did not have a large impact on performance. This is due to the timeout mechanism in place; if batch filling exceeds the timeout, the process is stopped, and the batch is processed with its current size.
(a) 2GPU in RC
(b) 2GPU in STP
(c) SingleGPU in RC
(d) SingleGPU in STP
Figure 3: Hyperparameter analysis for Batch IDA*.
@@
GPU Count & Hardware
To evaluate the impact of GPU count, we present the performance of Batch IDA* with different numbers of GPUs in Table 3, using another server for this experiment. In both domains, performance improves with additional GPUs, with the 4-GPU version matching AIDA*’s performance on the Rubik’s Cube. However, the performance gains diminish with increasing GPU counts, due to fewer threads being available to populate batches. Additionally, we assess Batch IDA* using various hardware configurations, with results provided in the Supplementary Material.
@@
We conclude by evaluating neural heuristics in Batch IDA* against AIDA*. Batch algorithms rely on admissible neural network heuristic $h_{M}$, while non-learning algorithms use a compressed PDB heuristic $h_{\text{DIV}}$ of equal size. Results for STP and Rubik’s Cube are presented in Table 4 and Table 5, respectively. The slower performance of learning algorithms compared to the results in Table 1 is due to the need for additional tensor operations per node to ensure heuristic admissibility. Non-learning algorithms expand more nodes in both domains since $h_{\text{DIV}}$ is a weaker heuristic than $h_{M}$. AIDA* performs significantly better than all other algorithms, while among the learning algorithms, Batch IDA* remains faster than Batch A* in terms of constant time per node. We provide a Detailed analysis of this performance gap between AIDA* and Batch IDA* in the Supplementary Material.
@@
This paper has shown how to parallelize CB-DFS on the CPU and GPU for efficiently using NN heuristics during search. CB-DFS is explored with Batch IDA*, but could also be in used in BTS if needed. The batching approach provides significant gains when batch sizes reach a sufficiently large size. Thus, this work provides the foundation for future research on building neural heuristics.
Acknowledgments
This work was supported by the National Science and Engineering Research Council of Canada Discovery Grant Program and the Canada CIFAR AI Chairs Program. This research was enabled in part by support provided by Prairies DRI and the Digital Research Alliance of Canada (alliance-can.ca). We extend deepest appreciation to our colleague who contributed significant feedback to improving this work.

### v2

@@ Introduction
Batch CB-DFS with Multiple GPUs
In most servers, multiple GPU devices are available, although the design shown thus far only incorporates a single GPU. If there are many fast CPUs available, they can end up waiting on the GPU for heuristic evaluations.
We scale Batch CB-DFS to multiple GPUs by loading $h_{M}$ onto each GPU, and running one batch-processing thread per GPU. CB-DFS threads are then uniformly assigned to the available GPUs, with each GPU processing the batch of nodes generated by its respective CB-DFS threads. This setup maximizes CPU utilization by increasing the number of batches while ensuring sufficient GPU resources are available to process them concurrently.
Correctness of Batch CB-DFS
In this section, IDA*, BTS, and AIDA* are all implemented with a CB-DFS sub-routine. Thus, as long as Batch CB-DFS searches the same tree as a classic CB-DFS search, we can substitute Batch CB-DFS without impacting the correctness of these algorithms. The batching operation does not impact the fact that the cost limit in the search is bounded, because the cost bound is still respected in the subtree expansions. Detailed proofs are found in supplementary material.
Batch BTS Enhancements
One important feature of the BTS algorithm is that it avoids the worst case where the IDA* tree grows too slowly, which can happen when there are many unique $h$-costs (e.g. as a result of using a regression-based heuristic). BTS guarantees that in each iteration the number of nodes will grow exponentially. It does this by establishing a limit of how fast the tree should grow in the next iteration (measured by node expansions), and then does a search over the cost bounds to find the cost bound that achieves this growth. Thus, in Batch BTS we must impose an additional node expansion limit in the CB-DFS. For efficiency, we do not check this after every expansion, but often enough to ensure that the search will terminate within a constant bound of the limit.
Next, we note that using classic heuristics, there is little cost for running many small iterations at the very beginning of search. But, with neural heuristics, doing so is very slow, because the search never grows large enough to exploit large batch sizes. Thus, a very small iteration of BTS can be a thousand times more expensive than a larger iteration.
To mitigate this, we make two additional modifications in Batch BTS. First, instead of setting the initial node limit to a small constant, such as 1, we use a much larger constant – in our case 10,000 expansions. This allows the initial searches to be significantly larger. But, setting the node limit on its own is not sufficient, we also need to raise the cost bound of the first search. Thus, instead of using $h(\text{root})$, we use $h(\text{root})+1$ instead. Neither of these changes impact the asymptotic behavior of the algorithm. Because the overheads are just large constants, on sufficiently large problems these modifications would not be necessary. But, in practice, they significantly improve performance.
@@
The memory cost of IDA* is $\mathcal{O}(d)$, where $d$ is the solution depth (korf1985depth). For AIDA*, which utilizes $n$ CB-DFS threads, the memory cost increases to $\mathcal{O}(dn)$ as it concurrently expands $n$ subtrees. In Batch IDA* and Batch BTS’s CB-DFS, each thread manages $k$ subtrees in a stack, resulting in a memory cost of $\mathcal{O}(dnk)$. We know that $n$ and $k$ are related to hardware speed on any domain. So, these parameters are expected to be constant as hardware is constant. Additionally, while the heuristic evaluations for unexpanded nodes are stored temporarily, memory is freed after their expansion, keeping the total memory cost at $\mathcal{O}(dnk)$.
AIDA*, Batch IDA*, and Batch BTS require loading their respective heuristics into memory. If AIDA* uses a PDB heuristic $h$ and Batch IDA* uses a learned model $h_{M}$, the total memory cost for AIDA* is $\mathcal{O}(nd+size(h))$ and for Batch IDA* it is $\mathcal{O}(dnk+size(h_{M}))$. While PDB heuristics typically fit into memory and thus their cost is constant, large PDBs, such as the 12-edge Rubik’s cube PDB (500 GB), exceed memory capacity, making them impractical for AIDA*. This limitation underscores the importance of further work on compressing PDB heuristics using deep learning and designing search algorithms, such as the CB-DFS used by Batch IDA* and Batch BTS, to use them efficiently.
@@
SingleGPU Batch IDA* 11 >> 100 - - $16.80$ $104,869$ $225,542$
SingleGPU Batch IDA* 88 >> 100 - - $3.97$ $93,610$ $201,217$
@@
Table 1: Summary results comparing Batch IDA* and Batch A* using the same PDB heuristic.
Experimental Results
The primary purpose of the experimental results is to evaluate whether or not the batched version of CB-DFS can effectively reduce the overhead of neural heuristics through GPU parallelism. We evaluate this in both the Batch IDA* and Batch BTS algorithms with neural heuristics built on both classifiers and regression, as regression-based heuristics are expected to fail with IDA*, but not BTS. We further evaluate the impact of GPU count and neural heuristic size on performance. Additional experiments examining the effects of hardware and algorithmic hyperparameters are included in the supplementary material.
Experimental Setup
We conduct our experiments across two domains: the 3x3 Rubik’s cube and the 4x4 sliding tile puzzle (STP). We use relatively modest computational resources in our experiments. Specifically, we utilize a server equipped with 32 AMD Ryzen Threadripper 2950X CPU cores and two NVIDIA GeForce RTX 2080 Ti GPUs running CUDA version 12.4. The Batch algorithms are implemented in C++, and Libtorch is employed for the deep learning components. Any CPU parallelization uses one thread per CPU core. For batched algorithms, we use a 4-millisecond timeout to trigger batch processing when an insufficient number of states are available to fully populate the batch. For the standard algorithms, IDA* and AIDA*, we use $h_{\text{PDB}}$ which is the 8-corners PDB for Rubik’s cube, and the sum of 1-7 and 8-15 tile additive PDBs for STP. For the STP, we use standard benchmark instances (korf1985depth). For the Rubik’s Cube domain, unless otherwise stated, we generated instances of solution length 11 using random walks starting from the goal state. PDB heuristics and baselines are from HOG2111https://github.com/nathansttt/hog2/tree/PDB-refactor.
The heuristic $h_{M}$ for the STP domain is the model introduced by li2022optimal. It is constructed by combining two ensemble models: the first is trained on the difference between the 1–7 tile PDB and the Manhattan distance heuristic, while the second is similarly trained using the 8–15 tile PDB. For the Rubik’s Cube domain, we trained two admissible heuristics $h_{M}$, both based on the 8-corner PDB, with a size of 1.9 MB. The first model is a classifier that was trained extensively to achieve both admissibility and the same average heuristic as in the input PDB. The second model is a regression network trained using mean squared error (MSE) loss. For admissibility, we subtracted 2.2 from all predicted heuristic values. However, this adjustment reduced the average heuristic by 2.3 compared to the perfect heuristic.
Fixed-Tree Evaluation
Our first experiment is designed to evaluate the effectiveness of batch operations in Batch IDA* with a classifier-based heuristic. We designed this experiment to isolate as many variables as possible in the experiment and to isolate the quality of neural heuristic from the mechanics of Batch IDA*. We isolate the impact of the neural heuristic training by fixing the tree to only contain states expanded by a baseline PDB heuristic ($h_{\text{PDB}}$). The search evaluates the neural heuristic at each state, but uses $h_{\text{PDB}}$ for pruning.
Under this methodology all algorithms are searching exactly the same tree. This allows us to initialize the neural heuristic randomly and quickly experiment with different network topologies without having to re-train the neural heuristics. This approach can measure the effectiveness of parallelism, but does not measure the quality of any learned heuristics, which is not the primary concern of this paper.
[…]
Average Performance Over 50 RC Instances
Algorithm Batch Size Time ± SE (s) Expanded
Solution Length 8
Batch BTS $10$ 33.41 ± 5.80 $537,763$
Batch BTS $100$ 4.26 ± 0.54 $569,475$
Batch BTS $1,000$ 0.97 ± 0.13 $571,179$
Batch BTS $2,000$ 0.83 ± 0.11 $583,135$
Batch IDA* $1,000$ 163.20 ± 54.63 $366,530,531$
Batch IDA* $2,000$ 144.91 ± 37.63 $360,097,779$
Solution Length 9
Batch BTS $10$ 169.73 ± 65.58 $1,211,327$
Batch BTS $100$ 20.25 ± 6.58 $1,217,398$
Batch BTS $1,000$ 3.64 ± 0.94 $1,294,972$
Batch BTS $2,000$ 3.01 ± 0.94 $1,294,329$
Batch IDA* $2,000$     -     -
Table 2: Performance on 3x3 Rubik’s Cube instances with the regression model.
@@
Regression-Based Neural Heuristics
In Table 2 we evaluate the impact of using a real-valued heuristic on Batch IDA* and Batch BTS. In this experiment both algorithms are using an identical neural heuristic, and we are measuring the cost of searching the resulting tree with that algorithm. Because the heuristic is learned using regression, which produces real-valued heuristics, we the IDA* iterations to grow very slowly. This is confirmed by the experimental results.
Due to the poor performance of Batch IDA*, we only report results where the algorithm is able to solve all instances within the 4-hour time limit. Batch IDA* expands orders of magnitude more states than Batch BTS, reinforcing that Batch BTS is better suited for regression-based heuristics. On instances with solution length 9, Batch IDA* expands billions of nodes per instance and fails complete within the time limit, while Batch BTS successfully solves all instances across all batch sizes. Notably, Batch BTS achieves a $56\times$ speedup when the batch size is increased from 10 to 2000.
@@
Our third experiment looks at the impact of model size on CB-DFS performance. We run Batch IDA* on the same domains with a single GPU and the best parameters from Table 1. We vary the model size by fixing the number of layers, but changing the number of fully connected neurons in each layer. The results of this experiment are found in Table 3. These results show that increasing the size of the model does not linearly increase the time required to evaluate the model, although larger models are more expensive to evaluate. Thus, there will be trade-offs between model quality and speed when deploying neural heuristics.
The results also show that Batch A* is less affected by model size than Batch IDA*. The main reason for this difference is that the batches in Batch IDA* are not always full, meaning the GPU isn’t being used as efficiently. This can be caused by node generations above the cost thresholds and the uneven distribution of work between subtrees. When the model requires more time to evaluate, these issues are accentuated. See the Supplementary Material for further analysis.
@@
GPU Count
To evaluate the impact of GPU count, we present the performance of Batch IDA* with different numbers of GPUs in Table 4, using other servers for this experiment. In both domains, performance improves with additional GPUs, with the 4-GPU version matching AIDA*’s performance on the Rubik’s Cube. However, the performance gains diminish with increasing GPU counts, due to fewer threads being available to populate batches. Additionally, we assess Batch IDA* using various hardware configurations, with results provided in the Supplementary Material.
@@
We conclude by evaluating neural heuristics in Batch IDA* against AIDA*. Batch algorithms rely on admissible neural network heuristic $h_{M}$, while standard algorithms use a compressed PDB heuristic $h_{\text{DIV}}$ of equal size. Results for STP and Rubik’s Cube are presented in Table 5 and Table 6, respectively. The slower performance of learning algorithms compared to the results in Table 1 is due to the need for additional tensor operations per node to ensure heuristic admissibility. Standard algorithms expand more nodes in both domains since $h_{\text{DIV}}$ is a weaker heuristic than $h_{M}$. AIDA* performs significantly better than all other algorithms, while among the batch algorithms, Batch IDA* remains faster than Batch A* in terms of constant time per node. We provide a Detailed analysis of this performance gap between AIDA* and Batch IDA* in the Supplementary Material.
@@
This paper has shown how to parallelize CB-DFS on the CPU and GPU for efficiently using neural heuristics during search. CB-DFS is explored with both Batch IDA and Batch BTS. The batching approach provides significant gains when batch sizes reach a sufficiently large size. Thus, this work provides the foundation for future research on building larger and stronger neural heuristics.
Acknowledgements
This work was supported by the National Science and Engi- neering Research Council of Canada Discovery Grant Pro- gram and the Canada CIFAR AI Chairs Program.
@@ Appendix A Appendix
Hyperparameter Analysis
[…]
(a) 2GPU in RC
(b) 2GPU in STP
(c) SingleGPU in RC
(d) SingleGPU in STP
Figure 4: Hyperparameter analysis for Batch IDA*.
Proofs of Theoretical Results
Lemma 1.
Given an admissible monotone cost function, iterative-deepening- A* will find a solution of least cost if one exists (korf1985depth).
Lemma 2.
Proof.
From Lemma 1, we know that IDA* using the cost function $f$ finds an optimal solution if a solution exists. Therefore, our task is to demonstrate that AIDA*, using the same cost function $f$ as IDA*, expands all the nodes that IDA* expands at each cost threshold. We know that, like IDA*, AIDA* does not increase the cost threshold until all nodes with costs below the current threshold have been expanded. Assume, for the sake of contradiction, that there exists a node $s_{missed}$ that IDA* expands but AIDA* does not. Let us also assume that the initial phase in AIDA* proceeds up to a depth of $d_{init}$. We select a path from the root to $s_{missed}$, denoted as pathpath. The node $s_{missed}$ falls into one of two categories:
  1. 1.
    The length of pathpath is less than or equal to $d_{init}$.
  2. 2.
    The length of pathpath exceeds $d_{init}$.
If $s_{missed}$ falls within the first category, it must be expanded during initial phase, as this phase generates the entire tree up to $d_{init}$ levels. Therefore, in this case, $s_{missed}$ cannot be missed by AIDA*. If $s_{missed}$ falls into the second category, we can express the pathpath as $\{root,s_{1},...s_{d_{init}},...,s_{missed}\}$. The pathpath intersects the tree generated in the initial phase at node $s_{d_{init}}$. We know that, during second and third phases, AIDA* grows the subtree for all leaf nodes of the initial tree, including $s_{d_{init}}$. Since both AIDA* and IDA* use the same cost function $f$, the same cost threshold, and both employ a depth-first search strategy, AIDA* is guaranteed to expand the node $s_{missed}$ during subtree expansion of $s_{d_{init}}$. Thus, AIDA* does not miss any node that IDA* expands. This concludes the proof that AIDA* will find an optimal solution if one exists. ∎
Theorem 1.
Proof.
  1. 1.
  2. 2.
    $h_{M}(s)\leq h(s)\ \ \ \ \forall s\in S.$